What is the acceleration due to gravity on the top of Mount Everest Mount Everest is the highest moun- tain peak of the world at the height of 8848 m The value at sea level is 9.80 m/sec2

Grade:11

1 Answers

Madabhushi
7 years ago

gh= gR2 / (R + h)2 = g / (1 + h/R )2 as h < < R it can be approximated as
gh= g ( 1 –2h/R) here radius of the earth is R = 6400Km = 6400 x 103m
So , on the Mount Everest the acceleration due to gravity is
gh= 9.80 ( 1 – 2 × 8848 / 6400000) = 9.77m/s2